OpenStudy (anonymous):

In any triangle ABC, if a=18,b=24,c=30,then find sinA, sinB and sinC.

OpenStudy (anonymous):

law of cosines then use law of sines

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[c^2=a^2+b^2-2ab \cos(C)\] \[\frac{\sin A}{a}=\frac{\sin B}{b}=\frac{\sin C}{c}\]

OpenStudy (anonymous):

the triangles sides given by u is a pythogarian triplet as 18^2 +24^2=324+576=900=30^2 thus <C =90 thus sinC=1 sinA= 18/30 =3/5 and sinB=24/30=4/5

OpenStudy (anonymous):

i want the sol. by law of sines and cosines

OpenStudy (anonymous):

@completeidiot continue your sol.

OpenStudy (anonymous):

nothing to complete the first equation is law of cosines, the 2nd is law of sines use those equations

OpenStudy (anonymous):

actually sinA/a=sinC/c bcoz C=90 hence sinA=a sinC /c=18*sin90/30=18/30=3/5 and same for sinB

OpenStudy (anonymous):

sinA=1/2 and sinB=2 and sin C=1/2

OpenStudy (anonymous):

AB=OB-OC=b-a same for BC and AC

OpenStudy (anonymous):

Then we get AB=6 BC=6 AC=12 sinA=6/12 and SinB=12/6 and sinC=6/12